A town has a population of 17000 and grows at 3% every year. What will be the population after 10 years, to the nearest whole number?
1
Expert's answer
2020-06-04T17:39:11-0400

Solution.

For the first year, the population of the city will increase by:
17000 + 17000 * 0,03 = 17510
For 2 year:
17510 + 17510 * 0,03 = 18035,3
For 3 year:
18035,3 + 18035,3 * 0,03 = 18576,359
For 4 year:
18576,359 + 18576,359 * 0,03 = 19133,6498
For 5 year:
19133,6498 + 19133,6498 * 0,03 = 19707,6593
For 6 year:
19707,6593 + 19707,6593 * 0,03 = 20298,8891
For 7 year: 
20298,8891 + 20298,8891 * 0,03 = 20907,8558
For 8 year:
20907,8858 + 20907,8858 * 0,03 = 21535,1224 
For 9 year:
21535,1224 + 21534,1224 * 0,03 = 22181,1761
For 10 year:
22181,1761 + 22181,1761 * 0,03 = 22846,6114 = 22847

Answer: in 10 years the population will be 22847 people